# Sparse Bayesian Multidimensional Item Response Theory

### v2

@@ 1 Introduction
The rest of this paper is organized as follows. Section 2 offers a concrete motivating example exhibiting the potential of our proposed approach to estimating sparse Bayesian MIRT models. Section 3 provides a detailed exposition of the data generation process, identification conditions, and the theoretical justifications behind our proposed hierarchical model. We then show the derivations of our parameter-expanded EM algorithm in Section 4. Section 5 further demonstrates the generality of our proposed methodology by accommodating mixed data types, including continuous and ordinal responses. In Section 6, the effectiveness of our proposed method is substantiated through its application to diverse scenarios: a challenging synthetic dataset, a large sample educational assessment dataset, a quality-of-life measurement dataset, and a challenging ordinal bio-behavioral dataset. We conclude the paper in Section 7 and provide additional details in the Appendix.
@@ 2 Exact Sparsity Recovery
Prior to delving into technical details, we provide a simple motivating example to illustrate the usefulness of our Bayesian approach when the factor loading matrix is sparse and its dimensionality uncertain. To this end, we generated four-factor, two-parameter binary item response data, and focused on a challenging high-dimensional $N<J$ case, in which $50\%$ of items load on an arbitrary pair of latent factors. As shown in the leftmost subplot in Figure 1, the true loading matrix is relatively sparse - the red areas represent zeros, while the yellow areas represent ones. Although simple looking, this loading structure can be very hard to recover in practice, as there exists a substantial overlap among latent factors.
@@ 3 Nonparametric Bayesian MIRT
The data generation process is finalized upon specifying priors for both the latent factor $\theta_{i}$ and the intercept $d_{j}$. In the context of exploratory factor analysis, it is customary to assume a prior of zero-mean multivariate Gaussian with an identity covariance matrix for the latent factors, since our goal is to unveil the latent structure of factor loadings. We further consider a Laplacian prior with small parameter $\lambda_{1}$ for the intercept parameter $d_{j}$. The combination of SSL and IBP priors on the loading matrix enhances the flexibility and interpretability of our model. This combination has exhibited notable success in promoting a small effective latent dimensionality and facilitating sparse estimation of the loading matrix within the context of linear factor models [bfa].
@@ 3.1 Probit Identifiability
Once we know the factor loading matrix with the GLT structure is uniquely identified, a natural question is whether we can achieve posterior consistency of $B$ under such a structural assumption. Notably, recent work by [Ma2020OnPC] advocates for a “$\sqrt{n}$-orthonormal factor model” approach under the SSL-IBP prior linear factor model setup, in which they consider the LQ decomposition of the $N$ by $K$ factor matrix $\Theta$ into the product of a lower triangular matrix $K(\Theta)$ and an orthogonal matrix $V(\Theta)$. After showing the consistency of $V(\Theta)$ as $J\to\infty$, they establish consistency of $B$ up to sign permutation as $N\to\infty$, condition on the feature allocation matrix $\Gamma$ possessing the GLT structure. The reparametrization of $\Theta$ help address the ”magnitude inflation problem” in high-dimensional setting, effectively mitigating sensitivity to the prior. In Appendix E.2, we empirically verify the consistency of $B$ with a challenging high-dimensional example as described in Section 6.2.
3.2 Factor Dimensionality
We provide robust theoretical justification for our methodology in recovering the true number of factors under the probit factor model. The effective dimension of the loading matrix $K(B)$ under the SSL-IBP($\{\lambda_{0k},\lambda_{1},C\}$) prior is defined as the largest column index $K$ such that $|B_{jk}|<\frac{1}{\lambda_{0k}-\lambda_{1}}\log[\frac{\lambda_{0k}(1-c_{k})}{\lambda_{1}c_{k}}]$ for any $j\in\{1,\cdots,J\}$ and $k>K$ [bfa]. Denote $B_{0n}\in\mathbbm{R}^{J_{n}\times\infty}$ as the true loading matrix with $K_{0n}$ nonzero columns, $\Theta_{0n}\in\mathbbm{R}^{n\times\infty}$ as the true latent factor matrix, and $\eta_{0ij}=B_{0j}^{\prime}\theta_{0i}$. For $1\leq k\leq K_{0n}$, let $S_{nk}:=\sum_{j=1}^{J_{n}}\mathbbm{I}(B_{jk}\neq 0)$ and $\bar{S}_{n}=\frac{1}{K_{0n}}\sum_{k=1}^{K_{0n}}S_{nk}$. We show the average posterior probability of $K(B)$ overshooting a multiple of the true sparsity level diminishes to zero asymptotically under the following assumptions: (A) $N<J$ and $N\log(\sqrt{N})<\log(J_{n}+1)K_{0n}\bar{S}_{n}$; (B) $\|B_{0n}\|_{2}<\sqrt{\bar{S}_{n}}$, where $\|B_{0n}\|_{2}$ is the spectral norm of the submatrix formed by the first $K_{0n}$ columns $\|B_{0n,K_{0n}}\|_{2}$; and (C) $\log\gamma_{n}(B_{0n},\Theta_{0n})\precsim\log J_{n}$, where $\phi$ is standard normal PDF, and
$\gamma_{n}(B_{0n},\Theta_{0n}):=1+\max_{\{i\leq n,j\leq J_{n}\}}\frac{\phi(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}.$
Assumptions A and B are identical to the assumptions made in [bfa], in which A restricts our analysis on the high-dimensional case, and B holds with high probability [bfa]. Assumption $C$ is an additional sparsity assumption on $(B_{0n},\Theta_{0n})$, analogous to a special case of Assumption 1 appeared in [glm_contraction] under the sparse generalized linear model setting.
@@ 3.2 Factor Dimensionality
Under assumptions (A)-(C), consider the sparse Bayesian MIRT model defined in equation (1), where the true loading matrix $B_{0n}$ has $K_{on}$ nonzero leftmost columns. For $1\leq k\leq K_{0n}$, let $\sum_{j=1}^{J_{n}}\mathbbm{I}(B_{jk}\neq 0)=S_{nk}$ and $\bar{S}_{n}=\frac{1}{K_{0n}}\sum_{k=1}^{K_{0n}}S_{nk}$. Then with the prior choice $B\sim\text{SSL-IBP}(\{\lambda_{0k}\},\lambda_{1},\alpha_{n})$, where $\alpha=\frac{1}{J_{n}}$, $\lambda_{0k}\geq\frac{2J_{n}^{2}k^{3}}{n}$, and $\lambda_{1}<e^{-2}$, and with some constant $C_{0}>0$:
$E_{B_{0}}P[K(B)>C_{0}K_{0n}\bar{S}_{n}|Y^{(n)}]\xrightarrow[n\rightarrow\infty]{}0.$
Theorem 3.4 provides a theoretical guarantee of our proposed Bayesian MIRT model in identifying the true latent factor dimensionality, a result analogous to Theorem 3 of [bfa] under a linear setting. Note the proof techniques of these two results are entirely different: in a linear setting, one may easily integrate Gaussian factor over the likelihood to study the posterior distribution of the covariance matrix under the SSL-IBP prior. This approach is no longer possible under a probit link. Hence we replace the assumptions on the true covariance matrix with assumption (C), which is used to control the Kullback-Leibler divergence and its variation between the posterior density and the true density. The corresponding proof can be found in Section B.2.
@@ 6.3 Massachusetts DESE Data
Figure 3 presents the estimated factor loading matrix for each model. We’ve labeled each row to indicate its corresponding item type as either English or math. Notably, our PXL-EM algorithm demonstrates a distinct pattern: high factor loadings for the secondary factor commence at item m22222222, aligning precisely with our expectations. Impressively, all math items exhibit substantial positive loadings on the second factor, yet the second factor for the English items are negligible, rendering our estimated loading matrix considerably more interpretable compared to the other models. In contrast, bboth the Gibbs sampler and the JML estimation algorithm managed to identify two prominent components, but neither of them successfully captured the distinction between the math and the english items. More importantly, for all the competing exploratory methods, their loading estimates for the extra dimensions $3$-$10$ significantly deviate from zero, making it very difficult to determine either the number of factors, or the potential latent structure.
@@ 6.4 Quality of Life Measurement
The loading matrix produced by our PXL-EM algorithm only differs from the oracle bifactor structure in two subdomains: the additional dimensionality for the “Health” items, and the absence of “Leisure” factor account for the FDR and FNR respectively as shown in Table 3. In contrast, both the MHRM algorithm and the Gibbs sampler utilized all $10$ dimensions, hence overfitting the item response data and having a marginally lower in-sample reconstruction MSE. The JML estimation method converged dramatically fast for this small $N$ dataset, but obtained relatively high estimation MSE, and a noisy loading representation. In contrast, the red areas visualized by our PXL-EM algorithm are exact zeros. We provide the detailed loading matrices and more discussions in Appendix G.
@@ 6.5 Ordinal bio-behavioral Data
As indicated by the color palette, the red, orange, and yellow regions represent negative, zero, and positive loadings respectively. Notably, the loading matrix estimated by PXL-EM is significantly sparser and more interpretable compared to those produced by the other methods: the first 1-3 dimensions represent the primary bio-behavioral dimensions of interest, shared by both the cortical thickness items (first 16 rows) and the clinical symptom items (last 51 rows). Dimension 4 is an independent factor for cortical thickness, while dimensions 5-10 are shared by the 51 clinical symptoms.
@@
To investigate the relationship between cortical thickness and clinical psychotic manifestations, it is essential to identify the (bio-behavioral) factors that are shared by all the items. Such analysis is unfeasible based on the factor loadings produced by the Gibbs sampler and the MHRM algorithm, where all the items are associated with all the $10$ factors, because these methods neither produce exact zeros nor can determine the number of factors. In contrast, our PXL-EM algorithm simplifies this investigation by identifying only three bio-behavioral dimensions, with the remaining seven dimensions independently loaded by either the biological or clinical features. More importantly, the second bio-behavioral dimension clearly highlights the negative correlation between cortical thickness and the severity of clinical symptoms, as the cortical thickness items are predominantly yellow, while the clinical symptoms items are overwhelmingly red.
@@ 7 Discussion
This article advocates for a novel Bayesian framework for estimating exploratory sparse MIRT models. The combination of the Indian Buffet Process prior and the spike-and-slab LASSO prior allows our EM algorithm to simultaneously learn latent dimensionality and sparse representations of factor loadings, which have not been sufficiently addressed in the binary and ordinal factor analysis literature. The efficiency of the algorithm is driven by the adaptation of parameter expansion, as well as by an efficient sampling strategy during the E-step, where we explicitly derived latent factor posterior distributions. We further demonstrate the advantage of adopting a probit link in establishing identification results and provide a theoretical guarantee for recovering the true number of factors under our prior specification. Our approach is capable of accommodating mixed data types, and we showcase the effectiveness of our approach on a challenging synthetic dataset and three real-world datasets from different domains. In each experiment presented in Section 6, our algorithm consistently generates sparse and interpretable estimations of the factor loading matrices, demonstrating its unique strength compared to other popular exploratory MIRT estimation frameworks.
@@ Appendix B Sparse Bayesian MIRT: Model Properties
$P(E|Y^{n})=\frac{\int_{E}\prod_{i=1}^{N}\prod_{j=1}^{J}f_{\pi}(y_{ij})/f_{\pi_{0}}(y_{ij})d\Pi}{\int\prod_{i=1}^{N}\prod_{j=1}^{J}f_{\pi}(y_{ij})/f_{\pi_{0}}(y_{ij})d\Pi}:=\frac{N_{n}}{D_{n}}.$
@@
$A_{n}:=\{B\in\mathbbm{R}^{J_{n}\times\infty},\Theta\in\mathbbm{R}^{N\times\infty}:\frac{1}{NJ_{n}}\sum_{i,j}K(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))\leq\frac{K_{0n}S_{n}\log(J_{n})}{NJ_{n}},$
$\frac{1}{NJ_{n}}\sum_{i,j}V_{2,0}(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))\leq\frac{K_{0n}S_{n}\log(J_{n})}{NJ_{n}}\}$
@@
$P_{\Pi_{0}}^{(n)}(D_{n}\geq\pi(A_{n})\cdot e^{-(1+c)K_{0n}S_{n}\log(J_{n})})\geq 1-\frac{1}{c^{2}K_{0n}S_{n}\log(J_{n})}\xrightarrow[]{n\to\infty}1$
@@
$K(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))=(\xi(\eta_{0_{ij}})-\xi(\eta_{{ij}}))b^{\prime}(\xi(\eta_{0_{ij}}))-b(\xi(\eta_{0_{ij}}))+b(\xi(\eta_{{ij}})),$
$V_{2,0}(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))=b^{{}^{\prime\prime}}(\xi(\eta_{0_{ij}}))(\xi(\eta_{0_{ij}})-\xi(\eta_{{ij}}))^{2}.$
@@
$\max\{K(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij})),V_{2,0}(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))\}$ (9)
$\quad\leq b^{{}^{\prime\prime}}(\xi(\eta_{0ij}))(\xi^{\prime}(\eta_{0ij}))^{2}(\eta_{ij}-\eta_{0ij})^{2}+O((\eta_{ij}-\eta_{0ij})^{2})$ (10)
$\quad=(h^{-1})^{{}^{\prime}}(\eta_{0ij})(\xi^{\prime}(\eta_{0ij}))(\eta_{ij}-\eta_{0ij})^{2}+O((\eta_{ij}-\eta_{0ij})^{2})$ (11)
$\quad=1+\max_{\{i\leq n,j\leq J_{n}\}}\frac{\phi(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}+O((\eta_{ij}-\eta_{0ij})^{2}).$ (12)
Recall $\gamma_{n}(B_{0n},\Theta_{0n}):=1+\max_{\{i\leq n,j\leq J_{n}\}}\frac{\phi(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}$. We can then lower bound $\pi(A_{n})$ as follows:
@@
To see why the last inequality is true for sufficiently large $n$, we can take $\log$ on both side. It is clear the second term $\frac{-nq_{n}}{2}\gtrsim-(K_{0n}S_{n}+1)(1+\log(J_{n}+1)$ due to the large denominator term. To see why the first term $-\frac{nK_{on}}{2}\log(2\pi/q_{n})\gtrsim-(K_{0n}S_{n}+1)(1+\log(J_{n}+1))$, we can divide the fractional term and then exponentiate both side, and it is sufficient to show $\frac{K_{0n}S_{n}\log(J_{n})}{2\pi\gamma_{n}n^{2}J_{n}(\sqrt{K_{0n}/n}+\sqrt{S_{n}J_{n}})^{2}}\geq(J_{n}+1)^{-c_{2}}$ for sufficiently large n and constant $c_{2}$. Since $\log(\gamma_{n})<\log(J_{n})$ by assumption C, $S_{n}>n$, and $N\log(\sqrt{N})<\log(J_{n}+1)K_{0n}S_{nk}$ by assumption A, we know the left hand side is lower bounded by the order of $J_{n}^{-4}$, and then with the choice $c_{2}>4$, the last inequality holds.

### v3

@@ 1 Introduction
The rest of this paper is organized as follows. Section 2 offers a concrete motivating example exhibiting the potential of our proposed approach to estimating sparse Bayesian MIRT models. Section 3 provides a detailed exposition of the data generation process, identification conditions, and the theoretical justifications behind our proposed hierarchical model. We then show the derivations of our parameter-expanded EM algorithm in Section 4. Section 5 further demonstrates the generality of our proposed methodology by accommodating mixed data types, including continuous and ordinal responses. In Section 6, the effectiveness of our proposed method is substantiated through its application to diverse scenarios: a challenging synthetic dataset, a large sample educational assessment dataset, a quality-of-life measurement dataset, and a challenging ordinal bio-behavioral dataset. We conclude the paper in Section 7 and provide additional details in the Appendix. The relevant code with tutorials can be found in the following Github repo: https://github.com/JiguangLi/sparse_bayesian_mirt.
@@ 2 Exact Sparsity Recovery
Prior to delving into technical details, we provide a simple motivating example to illustrate the usefulness of our Bayesian approach when the factor loading matrix is sparse and its dimensionality uncertain. To this end, we generated four-factor, two-parameter binary item response data, and focused on a challenging high-dimensional $N<J$ case, in which $50\%$ of items load on an arbitrary pair of latent factors. As shown in the leftmost subplot in Figure 1, the true loading matrix is relatively sparse - the light pink areas represent zeros, while the red areas represent ones. Although simple looking, this loading structure can be very hard to recover in practice, as there exists a substantial overlap among latent factors.
@@ 3 Nonparametric Bayesian MIRT
The data generation process is finalized upon specifying priors for both the latent factor $\theta_{i}$ and the intercept $d_{j}$. In the context of exploratory factor analysis, it is customary to assume a prior of zero-mean multivariate Gaussian with an identity covariance matrix for the latent factors, since our goal is to unveil the latent structure of factor loadings. We further consider a Laplacian prior with small parameter $\lambda_{1}$ for the intercept parameter $d_{j}$. The combination of SSL and IBP priors on the loading matrix enhances the flexibility and interpretability of our model, which also has exhibited notable success in the nonparametric linear factor model settings [bfa].
@@ 3.1 Probit Identifiability
Once we know the factor loading matrix with the GLT structure is uniquely identified, a natural question is whether we can achieve posterior consistency of $B$ under such a structural assumption. Notably, recent work by [Ma2020OnPC] advocates for a “$\sqrt{n}$-orthonormal factor model” approach under the SSL-IBP prior linear factor model setup, and establishes consistency of $B$ up to sign permutation, condition on the feature allocation matrix $\Gamma$ possessing the GLT structure. Their approach can also help address the ”magnitude inflation problem” in high-dimensional setting, effectively mitigating sensitivity to the prior.
3.2 Posterior Asymptotics
We provide robust theoretical justification for our proposed nonparametric MIRT model, demonstrating its capability to recover the true number of factors and achieving the desired posterior contraction rate for the factor loading matrix. The effective dimension of the loading matrix $K(B)$ under the SSL-IBP($\{\lambda_{0k},\lambda_{1},C\}$) prior is defined as the largest column index $K$ such that $|B_{jk}|<\frac{1}{\lambda_{0k}-\lambda_{1}}\log[\frac{\lambda_{0k}(1-c_{k})}{\lambda_{1}c_{k}}]$ for any $j\in\{1,\cdots,J\}$ and $k>K$ [bfa]. We denote $B_{0n}\in\mathbbm{R}^{J_{n}\times\infty}$ as the true loading matrix with $k_{0n}$ nonzero columns, $\Theta_{0n}\in\mathbbm{R}^{n\times\infty}$ as the latent factor matrix, and $\eta_{0ij}=B_{0j}^{\prime}\theta_{0i}$. For $1\leq k\leq k_{0n}$, let $S_{nk}=\sum_{j=1}^{J_{n}}\mathbbm{I}(B_{jk}\neq 0)$, $\bar{S}_{n}=\frac{1}{k_{0n}}\sum_{k=1}^{k_{0n}}S_{nk}$, and assume each column of $B_{0n}$ exhibits $S_{n}$-sparsity.
To study the asymptotic properties of the posterior, we make the following assumptions: (A) $n<J_{n}$ and $\lim_{n\to\infty}\sqrt{\frac{S_{n}k_{0n}\log^{2}J_{n}}{n}}=0$; (B) $\|B_{0n}\|_{2}<\sqrt{\bar{S}_{n}}$, where $\|B_{0n}\|_{2}$ is the spectral norm of the submatrix formed by the first $k_{0n}$ columns $\|B_{0n,k_{0n}}\|_{2}$; and (C) $\gamma_{n}(B_{0n},\Theta_{0n})\precsim\log J_{n}$, where $\phi$ is standard normal PDF, and
$\gamma_{n}(B_{0n},\Theta_{0n}):=1+\max_{\{i\leq n,j\leq J_{n}\}}\frac{\phi^{2}(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}.$
Assumptions A and B are similar to the assumptions made in [bfa], in which A restricts our analysis on the high-dimensional case, and B holds with high probability [bfa]. Assumption $C$ is an additional sparsity assumption on $(B_{0n},\Theta_{0n})$, analogous to a special case of assumption 1 appeared in [glm_contraction] under the sparse generalized linear model setting. Given that $\frac{\phi^{2}(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}$ grows linearly in $\eta_{0_{ij}}$ by the Mills ratio while $\max_{i,j}\eta_{0ij}$ roughly grows in the log scale, Assumption $C$ is relatively mild.
Theorem 3.4 shows the average posterior probability of our exploratory MIRT model overestimating the true number of factors, scaled by the true sparsity level, approaches zero asymptotically. This result is analogous to Theorem 3 in [bfa] within a linear setting.
@@ 3.2 Posterior Asymptotics
Under assumptions (A)-(C), consider the sparse Bayesian MIRT model defined in equation (1), where the true loading matrix $B_{0n}$ has $k_{0n}$ nonzero leftmost columns. For $1\leq k\leq k_{0n}$, let $\sum_{j=1}^{J_{n}}\mathbbm{I}(B_{jk}\neq 0)=S_{nk}$ and $\lambda_{0k}\geq\frac{2J_{n}^{2}k_{0n}^{3}n}{S_{nk}}$. Then with the prior choice $B\sim\text{SSL-IBP}(\{\lambda_{0k}\},\lambda_{1},\alpha_{n})$, where $\alpha=\frac{1}{J_{n}}$, $\lambda_{1}<e^{-2}$, and with some constant $C_{0}>0$:
$E_{B_{0}}P[K(B)>C_{0}k_{0n}S_{n}|Y^{(n)}]\xrightarrow[n\rightarrow\infty]{}0.$
The corresponding proof can be found in Section B.2. In the linear factor model setting [bfa], one may easily integrate Gaussian factor over the likelihood to study the posterior distribution of the covariance matrix. This approach is no longer possible under the probit link. Consequently, we introduce assumption (C) to control the Kullback-Leibler divergence and its variation between the posterior density and the true density.
[…]
Theorem 3.5.
Assuming the true loading matrix $B_{0n}$ has $k_{0n}$ columns, each of which is $S_{n}$-sparse. Under assumptions (A)-(C), and with the same SSL-IBP prior choice as in Theorem 3.4, these exists a constant $C_{0}^{\prime}$ such that
$E_{B_{0}}\Pi_{n}\left\{B:H_{n}(B,B_{0})>C_{0}^{\prime}\sqrt{\frac{k_{0n}S_{n}\log^{2}J_{n}}{n}}|Y^{(n)}\right\}\xrightarrow[n\rightarrow\infty]{}0.$
The Hellinger posterior contraction rate for the regression parameter in generalized linear models with a fixed design was recently shown to be $\sqrt{\frac{S_{0}\log J_{n}}{n}}$, where $S_{0}$ is the size of the true support [glm_contraction]. Theorem 3.5 provides similar contraction guarantee but with an extra $\sqrt{\log J_{n}}$ factor, reflecting the cost of the growing number of factors $k_{0n}$ and the support size $S_{n}$. The proof of Theorem 3.5 is detailed in Appendix B.3. In Appendix E.2, we further empirically verify the consistency of $B$ with a challenging high-dimensional example as described in Section 6.2.
@@ 6.3 Massachusetts DESE Data
Figure 3 presents the estimated factor loading matrix for each model. We’ve labeled each row to indicate its corresponding item type as either English or math. Notably, our PXL-EM algorithm demonstrates a distinct pattern: high factor loadings for the secondary factor commence at item m22222222, aligning precisely with our expectations. Impressively, all math items exhibit substantial positive loadings on the second factor, yet the second factor for the English items are negligible, rendering our estimated loading matrix considerably more interpretable compared to the other models. In contrast, both the Gibbs sampler and the JML estimation algorithm managed to identify two prominent components, but neither of them successfully captured the distinction between the math and the english items. More importantly, for all the competing exploratory methods, their loading estimates for the extra dimensions $3$-$10$ significantly deviate from zero, making it very difficult to determine either the number of factors, or the potential latent structure.
@@ 6.4 Quality of Life Measurement
The loading matrix produced by our PXL-EM algorithm only differs from the oracle bifactor structure in two subdomains: the additional dimensionality for the “Health” items, and the absence of “Leisure” factor account for the FDR and FNR respectively as shown in Table 3. In contrast, both the MHRM algorithm and the Gibbs sampler utilized all $10$ dimensions, hence overfitting the item response data and having a marginally lower in-sample reconstruction MSE. The JML estimation method converged dramatically fast for this small $N$ dataset, but obtained relatively high estimation MSE, and a noisy loading representation. In contrast, the pink areas visualized by our PXL-EM algorithm are exact zeros. We provide the detailed loading matrices and more discussions in Appendix G.
@@ 6.5 Ordinal bio-behavioral Data
As indicated by the color palette, the red, light pink, and blue regions represent negative, zero, and positive loadings respectively. Notably, the loading matrix estimated by PXL-EM is significantly sparser and more interpretable compared to those produced by the other methods: the first 1-3 dimensions represent the primary bio-behavioral dimensions of interest, shared by both the cortical thickness items (first 16 rows) and the clinical symptom items (last 51 rows). Dimension 4 is an independent factor for cortical thickness, while dimensions 5-10 are shared by the 51 clinical symptoms.
@@
To investigate the relationship between cortical thickness and clinical psychotic manifestations, it is essential to identify the (bio-behavioral) factors that are shared by all the items. Such analysis is unfeasible based on the factor loadings produced by the Gibbs sampler and the MHRM algorithm, where all the items are associated with all the $10$ factors, because these methods neither produce exact zeros nor can determine the number of factors. In contrast, our PXL-EM algorithm simplifies this investigation by identifying only three bio-behavioral dimensions, with the remaining seven dimensions independently loaded by either the biological or clinical features. More importantly, the second bio-behavioral dimension clearly highlights the negative correlation between cortical thickness and the severity of clinical symptoms, as the cortical thickness items are predominantly red, while the clinical symptoms items are overwhelmingly blue.
@@ 7 Discussion
This article advocates for a novel Bayesian framework for estimating exploratory sparse MIRT models. The combination of the Indian Buffet Process prior and the spike-and-slab LASSO prior allows our EM algorithm to simultaneously learn latent dimensionality and sparse representations of factor loadings, which have not been sufficiently addressed in the binary and ordinal factor analysis literature. The efficiency of the algorithm is driven by the adaptation of parameter expansion, as well as by an efficient sampling strategy during the E-step, where we explicitly derived latent factor posterior distributions. We further demonstrate the advantage of adopting a probit link for establishing identification results and provide theoretical guarantees for recovering the true number of factors and contraction towards the true factor loading matrix under our prior specification. Our approach is capable of accommodating mixed data types, and we showcase the effectiveness of our approach on a challenging synthetic dataset and three real-world datasets from different domains. In each experiment presented in Section 6, our algorithm consistently generates sparse and interpretable estimations of the factor loading matrices, demonstrating its unique strength compared to other popular exploratory MIRT estimation frameworks.
@@ Appendix B Sparse Bayesian MIRT: Model Properties
$P(E|Y^{n})=\frac{\int_{E}\prod_{i=1}^{N}\prod_{j=1}^{J}f_{\pi}(y_{ij})/f_{\pi_{0}}(y_{ij})d\Pi}{\int\prod_{i=1}^{N}\prod_{j=1}^{J}f_{\pi}(y_{ij})/f_{\pi_{0}}(y_{ij})d\Pi}:=\frac{N_{n}}{D_{n}}.$ (8)
@@
$A_{n}:=\{B\in\mathbbm{R}^{J_{n}\times\infty},\Theta\in\mathbbm{R}^{N\times\infty}:\frac{1}{NJ_{n}}\sum_{i,j}K(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))\leq\frac{K_{0n}S_{n}\log(J_{n})}{NJ_{n}},$ (9)
$\frac{1}{NJ_{n}}\sum_{i,j}V_{2,0}(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))\leq\frac{K_{0n}S_{n}\log(J_{n})}{NJ_{n}}\}$ (10)
@@
$P_{\Pi_{0}}^{(n)}(D_{n}\geq\pi(A_{n})\cdot e^{-(1+c)K_{0n}S_{n}\log(J_{n})})\geq 1-\frac{1}{c^{2}K_{0n}S_{n}\log(J_{n})}\xrightarrow[]{n\to\infty}1$ (11)
@@
$K(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))=(\xi(\eta_{0_{ij}})-\xi(\eta_{{ij}}))b^{\prime}(\xi(\eta_{0_{ij}}))-b(\xi(\eta_{0_{ij}}))+b(\xi(\eta_{{ij}})),$ (13)
$V_{2,0}(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))=b^{{}^{\prime\prime}}(\xi(\eta_{0_{ij}}))(\xi(\eta_{0_{ij}})-\xi(\eta_{{ij}}))^{2}.$ (14)
@@
$\max\{K(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij})),V_{2,0}(f_{\pi_{0}}(y_{ij}),f_{\pi}(y_{ij}))\}$
$\quad\leq b^{{}^{\prime\prime}}(\xi(\eta_{0ij}))(\xi^{\prime}(\eta_{0ij}))^{2}(\eta_{ij}-\eta_{0ij})^{2}+O((\eta_{ij}-\eta_{0ij})^{2})$
$\quad=(h^{-1})^{{}^{\prime}}(\eta_{0ij})(\xi^{\prime}(\eta_{0ij}))(\eta_{ij}-\eta_{0ij})^{2}+O((\eta_{ij}-\eta_{0ij})^{2})$
$\quad=1+\max_{\{i\leq n,j\leq J_{n}\}}\frac{\phi^{2}(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}+O((\eta_{ij}-\eta_{0ij})^{2}).$ (15)
Recall $\gamma_{n}(B_{0n},\Theta_{0n}):=1+\max_{\{i\leq n,j\leq J_{n}\}}\frac{\phi^{2}(\eta_{0_{ij}})}{\Phi(\eta_{0ij})(1-\Phi(\eta_{0ij}))}$. We can then lower bound $\pi(A_{n})$ as follows:
@@
To see why the last inequality is true for sufficiently large $n$, we can take $\log$ on both side. It is clear the second term $\frac{-nq_{n}}{2}\gtrsim-(K_{0n}S_{n}+1)(1+\log(J_{n}+1)$ due to the large denominator term. To see why the first term $-\frac{nK_{on}}{2}\log(2\pi/q_{n})\gtrsim-(K_{0n}S_{n}+1)(1+\log(J_{n}+1))$, we can divide the fractional term and then exponentiate both side, and it is sufficient to show $\frac{K_{0n}S_{n}\log(J_{n})}{2\pi\gamma_{n}n^{2}J_{n}(\sqrt{K_{0n}/n}+\sqrt{S_{n}J_{n}})^{2}}\geq(J_{n}+1)^{-c_{2}}$ for sufficiently large n and constant $c_{2}$. Since $\gamma_{n}\lesssim\log(J_{n})$ by assumption C, $J_{n}>n$ by assumption A, we know the left hand side is lower bounded by the order of $J_{n}^{-4}$, and then with the choice $c_{2}>4$, the last inequality holds.